OpenStudy (anonymous):

The length of a rectangle is 6 feet longer than its width. Find the dimensions of the rectangle assuming that its perimeter is 40 feet Width= Length=

OpenStudy (gorv):

parameter=2*(length+width) length=width+6 40/2=width+6+width 20-6=2 * width 14/2=width width=7 length=7+6=13

OpenStudy (gorv):

@TajaRice12

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

yes ????

OpenStudy (anonymous):

thanks @gorv

OpenStudy (gorv):

alwys welcome :)